show that f°f=2f

f=x+1


1
Expert's answer
2021-02-24T06:58:54-0500

"f(x)=x+1". The composition "f\circ f" has the form: "f\circ f(x)=f(x+1)=x+2";

"2f(x)=2x+2". As we can see, "f\circ f(x)\neq2f(x)".